OpenStudy (anonymous):

if you are graphing a function F(x)=x^2 what would f(x+4) would the graph be moving to the right 4 or to the left 4? help please

OpenStudy (anonymous):

niether left 4 nor right for... it would be a new parabola...i think

OpenStudy (anonymous):

i have four different graphs to choose from the other three choices besides those two are flipped from the top of the graph to the bottom on the left side

OpenStudy (anonymous):

x^2 +16+8x

OpenStudy (anonymous):

other 2 i meant

OpenStudy (anonymous):

so it moves to the left 4

OpenStudy (anonymous):

as of your equation unless u flip it or something

OpenStudy (amistre64):

(x+k) "moves" the original graph by -k

OpenStudy (amistre64):

just watch the difference in the vertex

OpenStudy (amistre64):

y = x^2, vertex is 0,0 y = (x+4)^2, y = 0 when x=-4, vertex is (-4,0)

OpenStudy (anonymous):



OpenStudy (anonymous):

this looks something like the orginal before that give u f(x+4)

OpenStudy (anonymous):

would it end up being like this